frank is 15 years younger than john.in 5 years john will be twice as old as frank,how old will frank be in 4 yeras

Answer by lwsshak3(11628)

frank is 15 years younger than john.in 5 years john will be twice as old as frank,how old will frank be in 4 years
***
let x=john"s age
x-15=frank's age
in 5 years:
x+5=john's age
x-15+5=x-10=frank's age
x+5=2(x-10)
x+5=2x-20
x=25
in 4 yrs
x-15+4=x-11=25-11=14
how old will frank be in 4 years=14